# Practical multi-fidelity machine learning: fusion of deterministic and Bayesian models

### v1

@@ 6 Conclusions
In this study, we created a general multi-fidelity framework that encompasses different multi-fidelity strategies in the literature. We propose using non-probabilistic LF models together with Bayesian HF models, and recommend two practical strategies to deal with data-scarce and data-rich scenarios. For the first case, we propose the the KRR-LR-GPR multi-fidelity model that considers kernel ridge regression for LF, linear regression for transfer learning data from LF to HF, and Gaussian process regression for modeling the residual that remains. For data-rich and high-dimensional cases, we recommend the use of DNN-LR-BNN that considers deep neural network for LF, linear regression transfer-learning, and Bayesian neural network for the residual. In case a linear regression model does not provide an effective transfer learning method, a more flexible DNN-BNN model is recommended using pSGLD as inference method for high-dimensional datasets.
The KRR-LR-GPR model is assessed for cases with and without noise for both fidelities and its performance is similar to the best state-of-the-art multi-fidelity GPR approaches, while having significantly shorter inference time (the KRR-LR-GPR method becomes bounded only by the number of HF samples). Finally, an ablation study shows the performance of KRR-LR-GPR is influenced by factors similar to those of other multi-fidelity GPR methods, so the observations align with those discussed in [8] for those methods.
We also report the performance of the DNN-LR-BNN compared with that of a single-fidelity BNN and a flexible multi-fidelity model considering only DNN-BNN. We observed that DNN-LR-BNN has lower tendency to overfit than DNN-BNN, as the LR transfer-learning model isolates the HF from the LF when they are not linearly correlated. Furthermore, we highlight the proposed strategy to determine $\bm{\rho}$ from training, instead of considering them as hyperparameters as in the literature.
Although the DNN-LR-BNN model can converge to the results of the KRR-LR-BNN model for cases when both are applicable, the KRR-LR-GPR model is easier to train because it has significantly fewer hyperparameters, hence the motivation of recommending it for data-scarce or lower dimensional scenarios.
@@
The code is available on: https:/https://github.com/bessagroup/mfbml
@@
The first author used CHATGPT and GRAMMARLY to IMPROVE LANGUAGE AND READABILITY. After using these tools/services, the authors reviewed and edited the content as needed and assume full responsibility for the content of the publication.
@@ Acknowledgments
Jiaxiang Yi would like to acknowledge the generous support of the China Scholarship Council (CSC).
@@ A.1 Bayesian neural networks
Consider first the nonprobabilistic DNNs. Fig.11a shows a DNN schematic, where the input $\mathbf{x}$ passes through successive neurons located in hidden layers where nonlinear transformations $\phi(\mathbf{x})$ happen in so called activation functions. In the end, the prediction $\hat{\mathbf{y}}$ is outputted by the DNN and compared to the true data via loss function. The solid lines that connect neurons include parameters called weights $\mathbf{w}$ while biases are not explicitly included in the schematic (together the weights and biases parameters are labeled as $\boldsymbol{\theta}$). These parameters are real values that are determined via optimization algorithms like Adam [63] and Stochastic Gradient Decent (SGD) [64] that minimize the loss function.
@@
BNNs are similar to DNNs, as shown in Fig.11b, but have a key difference: distributions over weights are considered instead of using a deterministic value as in DNNs. This probabilistic treatment allows BNNs to capture uncertainty in the predictions. Consequently, traditional optimization algorithms, suitable for finding point estimates, are not directly applicable to BNNs. The procedure to find the posterior parameter distribution is called inference, such that the posterior is determined as follows [47]:
@@
where $p\left(\boldsymbol{\theta}\right)$ is the prior, for example a normal distribution with $\boldsymbol{\theta}$ containing the mean $\boldsymbol{\mu_{\mathbf{w}}}$ and standard deviation of the weights $\boldsymbol{\sigma}_{\mathbf{w}}$; $p\left(\mathcal{D}\mid\boldsymbol{\theta}\right)$ is the likelihood; $p\left(\mathcal{D}\right)$ is the marginal likelihood or evidence, which is the probability integral of likelihood over $\boldsymbol{\theta}$; and $p\left(\boldsymbol{\theta}\mid\mathcal{D}\right)$ is the posterior distribution of parameter $\boldsymbol{\theta}$. The aim when training BNNs is to maximize $p\left(\boldsymbol{\theta}\mid\mathcal{D}\right)$, although this is usually intractable because of the unbounded and complex probability integral. Therefore, methods such as Markov Chain Monte Carlo [47] and Variational Inference [49] are extensively used to determine an approximate posterior distribution. After obtaining parameters, the prediction for an unknown point is calculated by:
@@ A.1.1 Bayesian inference method: pSGLD
As the analytical solution of the posterior predictive distribution of BNNs is intractable, we used preconditioned stochastic gradient Langevin dynamics (pSGLD) [58]. The posterior of a BNN can be regarded as $p(\bm{\theta}|\mathcal{D})\propto p(\bm{\theta})\prod_{n=1}^{N}p(y_{i}|\bm{\theta})$ according to Eq. (10), and we can define the update rule for the posterior as in SGLD [48]:
@@
where $\bm{\eta}_{t}\sim\mathcal{N}(\bm{0},\bm{\epsilon}_{t})$. However, SGLD assumes that all parameters $\bm{\theta}$ have the same step size, which can lead to slow convergence or even divergence in the cases where the components of $\bm{\theta}$ have different curvature. In pSGLD, the update rule incorporates a user-defined preconditioning matrix $G(\bm{\theta}_{t})$, which adjusts the gradient updates and the noise term adaptively:
@@
We can then obtain the posterior approximately with a set of $\bm{\Theta}=[\bm{\theta}_{1},\bm{\theta}_{2},...,\bm{\theta}_{n}]$ after $n$ updates from pSGLD. Then, the approximation of Eq. (11) can be obtained by a Bayesian model average [65].
@@ A.2.1 Gaussian process regression with explicit basis function
where $\mathbf{h}(\mathbf{x})$ is a set of basis function, i.e. $\mathbf{h}(\mathbf{x})=[1,\mathbf{x},\mathbf{x}^{2},...]^{T}$ where the length is set to be $M$; $\delta(\mathbf{x})$ is a GPR with zero mean expressed as $\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}))$, $\bm{\beta}$ is a set of additional parameters that used to tune basis function for minimizing the residual between $\mathbf{h}(\mathbf{x})^{T}\bm{\beta}$ and $\mathbf{y}$ (see Fig. 1). The parameters within the covariance function and $\bm{\beta}$ are determined by maximum likelihood estimation referred to Appendix. A.2.2. The predictions [^3] For simplification, Eqs. (15) and (16) are derived assuming that $\bm{\beta}$ is a deterministic value rather than another distribution. Readers can find more information in [14] for cases where $\bm{\beta}$ also follows a distribution. of the general GPR considering the effect brought by the explicit basis function can then be expressed as [^4] Eqs. (15) and (16) assume no noise in observations. If the response has noise that is independent identically distributed and Gaussian $\varepsilon$ with variance $\sigma^{2}$, the kernel function should have one extra term to characterize noise. To this end, the noise kernel function can be written as $\mathbf{K}_{noisy}\left(\mathbf{X},\mathbf{X}\right)=\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma^{2}\mathbf{I}$  [^5] Naming convention: $k(\mathbf{x},\mathbf{x})$ represents a number; $\mathbf{k}(\mathbf{x},\mathbf{X})$ represents a $1\times n$ vector and vice verse; $\mathbf{K}(\mathbf{X},\mathbf{X})$ is a $n\times n$ matrix. Moreover, the default vector is a column vector.:.
$f(\mathbf{x}^{\prime})=\mathbf{h}(\mathbf{x^{\prime}})^{T}\mathbf{\hat{\boldsymbol{\beta}}}+\mathbf{k}(\mathbf{x}^{\prime},\mathbf{X})\mathbf{K}(\mathbf{X},\mathbf{X})^{-1}(\mathbf{y}-\mathbf{h}(\mathbf{X})^{T}\mathbf{\hat{\boldsymbol{\beta}}})$ (15)
$s_{f}^{2}(\mathbf{x}^{\prime})=k(\mathbf{x}^{\prime},\mathbf{x}^{\prime})-\mathbf{k}\left(\mathbf{x}^{\prime},\mathbf{X}\right)\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)^{-1}\mathbf{k}\left(\mathbf{X},\mathbf{x}^{\prime}\right)+\mathbf{r}^{T}(\mathbf{h}(\mathbf{X})\mathbf{K}(\mathbf{X},\mathbf{X})^{-1}\mathbf{h}(\mathbf{X})^{T})^{-1}\mathbf{r}$ (16)
where $\mathbf{r}=\mathbf{h}(\mathbf{x^{\prime}})-\mathbf{h}(\mathbf{X})\mathbf{K}(\mathbf{X},\mathbf{X})^{-1}\mathbf{k}(\mathbf{X},\mathbf{x^{\prime}})$. The first two terms are predictive variance deduced by the zero mean GPR and the third term is the predictive variance caused by the explicit basis function.
Incorporating explicit basis functions helps the GPR to generalize for more complicated problems, while one has to select the basis function $\mathbf{h(x)}$ beforehand or use symbolic regression to identify the terms [66].
@@ A.2.2 Parameters estimation
By taking derivatives of Eq. (18) and setting to zero, we obtain maximum likelihood estimation for $\boldsymbol{\beta}$.
@@
Now the maximum likelihood estimations $\hat{\boldsymbol{\beta}}$ can be substituted back into Eq. (18) and constant terms removed to give the concentrated ln-likelihood function:
@@
The value of Eq. (20) is determined by parameter $\bm{\theta}$, the maximum value can not be found analytically. Therefore, heuristic algorithms can be utilized to find the optimum value of $\bm{\theta}$.
@@ B.1 LF-KRR parameter estimation
Kernel ridge regression based on the LF dataset can be defined by
@@ B.2 Transfer-learning model and residual GPR parameter estimation
By taking the derivative of Eq. (23) and setting it to zero, we can get the estimation of $\hat{\bm{\rho}}$ :
@@ B.3 Predictions
We can obtain the predictions of KRR-LR-GPR by modifying Eqs. 14 and 16 into:
$f^{h}(\mathbf{x}^{\prime})=\mathbf{m}(\mathbf{x^{\prime}})^{T}\mathbf{\hat{\bm{\rho}}}+\mathbf{k}(\mathbf{x}^{\prime},\mathbf{X}^{h})\mathbf{K}(\mathbf{X}^{h},\mathbf{X}^{h})^{-1}(\mathbf{y}^{h}-\mathbf{m}(\mathbf{X}^{h})^{T}\mathbf{\hat{\bm{\rho}}})$ (26)
$s_{f^{h}}^{2}(\mathbf{x}^{\prime})=k(\mathbf{x}^{\prime},\mathbf{x}^{\prime})-\mathbf{k}\left(\mathbf{x}^{\prime},\mathbf{X}^{h}\right)\mathbf{K}\left(\mathbf{X}^{h},\mathbf{X}^{h}\right)^{-1}\mathbf{k}\left(\mathbf{X}^{h},\mathbf{x}^{\prime}\right)+\mathbf{r}^{T}(\mathbf{m}(\mathbf{X}^{h})\mathbf{K}(\mathbf{X}^{h},\mathbf{X}^{h})^{-1}\mathbf{m}(\mathbf{X}^{h})^{T})^{-1}\mathbf{r}$ (27)
@@ C.1 Numerical functions for testing KRR-LR-GPR
This section includes details on the low-dimensional numerical functions considered in this paper. Originally, those functions were single fidelity functions used to verify optimization algorithms [67]. They were extended by Jiang et al [68], Sander et al. [69] and Mainini et al. [70] for verifying the performance of MF optimization algorithms.
@@
    where $x\in[0,1]$. By selecting different values of $A$, $B$, and $C$, LF functions with different correlations to the HF function. Three different configurations are adopted in Section 4.1 where $lf1$ considers $A=1,B=0,C=5$; $lf2$ considers $A=0.5,B=10,C=5$; $lf3$ considers $A=0.1,B=10,C=0.1$.

### v2

@@ 6 Conclusions
In this study, we created a general MF framework that encompasses different MF strategies in the literature. We propose using non-probabilistic LF models together with Bayesian HF models and recommend two practical strategies to deal with data-scarce and data-rich scenarios. For the first case, we propose the KRR-LR-GPR MF model that considers KRR for training on LF data, linear regression for LF to HF data transfer-learning, and Gaussian process regression for modeling the residual that remains. For data-rich, high-dimensional cases and/or multiple outputs, we recommend the DNN-LR-BNN model that involves training a deep neural network on LF data, transfer-learning via linear regression, and Bayesian neural networks for learning the residual. If a linear regression model does not provide an effective transfer-learning method, a more flexible DNN-BNN model is recommended using pSGLD for Bayesian inference.
The KRR-LR-GPR model was assessed for cases with and without noise for both fidelity levels, and its performance was found to be similar to state-of-the-art MF-GPR methods. However, KRR-LR-GPR exhibits two key advantages: 1) it can be trained on larger LF datasets, and 2) it is significantly faster to train (in practice, the method is only bounded by the number of HF samples, due to the cubic complexity of GPR models with the number of training samples). The method was also successfully applied to a dataset from an engineering application – prediction of aerodynamic coefficients of NACA 0012 airfoil. We believe that the simplicity of the training strategy and robustness of the method makes it suitable for most engineering applications because the most common scenario involves data-scarce HF samples (usually originating from expensive or time-consuming experiments or simulations).
We also proposed the DNN-LR-BNN model and compared it with both a single-fidelity BNN and a flexible MF model composed of a DNN and a BNN (DNN-BNN model). We observed that the DNN-LR-BNN model has a lower tendency to overfit when compared to DNN-BNN because its linear regression (LR) transfer-learning model isolates the HF from the LF when they are not linearly correlated. Furthermore, we highlight that the proposed strategy determines the coefficients of the LR model, $\bm{\rho}$, directly from training (instead of considering them as hyperparameters, as done in the literature for other models). The learned $\hat{\bm{\rho}}$ also indicates the correlation between HF and LF data, as shown in different examples such as the material structure-property relationship problem. Nevertheless, the DNN-LR-BNN model has more hyperparameters than the KRR-LR-GPR model, so it should be selected only when training the latter is not viable.
@@
The code is available on: https://github.com/bessagroup/mfbml
@@
The first author used CHATGPT and GRAMMARLY to IMPROVE LANGUAGE AND READABILITY. After using these tools/services, the authors reviewed and edited the content as needed and assumed full responsibility for the content of the publication.
@@ Acknowledgments
Jiaxiang Yi would like to acknowledge the generous support of the China Scholarship Council (CSC). We also appreciate Dr. Quan Lin for his kindness in providing the source data for the NACA 0012 airfoil problem.
@@ A.1 Bayesian neural networks
Consider first the nonprobabilistic DNNs. Figure 14.(a) shows a DNN schematic, where the input $\mathbf{x}$ passes through successive neurons located in hidden layers where nonlinear transformations $\phi(\mathbf{x})$ happen in so-called activation functions. To the end, the prediction $\hat{\mathbf{y}}$ is outputted by the DNN and compared to the true data via the loss function. The solid lines that connect neurons include parameters called weights $\mathbf{w}$ while biases are not explicitly included in the schematic (together the weights and biases parameters are labeled as $\boldsymbol{\theta}$). These parameters are real values that are determined via optimization algorithms like Adam [70] and Stochastic Gradient Decent [71] that minimize the loss function.
@@
BNNs are similar to DNNs, as shown in Figure 14.(b), but have a key difference: distributions acting on weights are considered instead of using a deterministic value as in DNNs. This probabilistic treatment allows BNNs to capture uncertainty in the predictions. Consequently, traditional optimization algorithms, suitable for finding point estimates, are not directly applicable to BNNs. The procedure to find the posterior parameter distribution is called inference, such that the posterior is determined as follows [46]:
@@
where $p\left(\boldsymbol{\theta}\right)$ is the prior, for example, a normal distribution with $\boldsymbol{\theta}$ containing the mean $\boldsymbol{\mu_{\mathbf{w}}}$ and standard deviation of the weights $\boldsymbol{\sigma}_{\mathbf{w}}$; $p\left(\mathcal{D}\mid\boldsymbol{\theta}\right)$ is the likelihood; $p\left(\mathcal{D}\right)$ is the marginal likelihood or evidence, which is the probability integral of likelihood over $\boldsymbol{\theta}$; and $p\left(\boldsymbol{\theta}\mid\mathcal{D}\right)$ is the posterior distribution of parameter $\boldsymbol{\theta}$. The aim when training BNNs is to maximize $p\left(\boldsymbol{\theta}\mid\mathcal{D}\right)$, although this is usually intractable because of the unbounded and complex probability integral. Therefore, methods such as Markov Chain Monte Carlo [46] and Variational Inference [48] are extensively used to determine an approximate posterior distribution. After obtaining parameters, the prediction for an unknown point is calculated by:
@@ A.1.1 Bayesian inference method: pSGLD
As the analytical solution of the posterior predicted distribution of BNNs is intractable, we used pSGLD [58]. The posterior of a BNN can be regarded as $p(\bm{\theta}|\mathcal{D})\propto p(\bm{\theta})\prod_{n=1}^{N}p(y_{i}|\bm{\theta})$ according to Equation 8, and we can define the update rule for the posterior as in SGLD [47]:
@@
where $\bm{\eta}_{t}\sim\mathcal{N}(\bm{0},\bm{\epsilon}_{t})$. However, SGLD assumes that all parameters $\bm{\theta}$ have the same step size, which leads to slow convergence or even divergence in the cases where the components of $\bm{\theta}$ have different curvature. In pSGLD, the update rule incorporates a user-defined preconditioning matrix $G(\bm{\theta}_{t})$, which adjusts the gradient updates and the noise term adaptively:
@@
We can then obtain the posterior approximately with a set of $\bm{\Theta}=[\bm{\theta}_{1},\bm{\theta}_{2},...,\bm{\theta}_{n}]$ after $n$ updates from pSGLD. Then, the approximation of Equation 9 can be obtained by a Bayesian model average [72].
@@ A.2.1 Gaussian process regression with explicit basis function
where $\mathbf{h}(\mathbf{x})$ is a set of basis function, i.e. $\mathbf{h}(\mathbf{x})=[1,\mathbf{x},\mathbf{x}^{2},...]^{T}$ where the length is set to be $M$; $\delta(\mathbf{x})$ is a GPR with zero mean expressed as $\mathcal{GP}(0,k(\mathbf{x},\mathbf{x}))$, $\bm{\beta}$ is a set of additional parameters that used to tune basis function for minimizing the residual between $\mathbf{h}(\mathbf{x})^{T}\bm{\beta}$ and $\mathbf{y}$ (see Figure 1). The parameters within the covariance function and $\bm{\beta}$ are determined by maximum likelihood estimation referred to Section A.2.2. The predictions [^3] For simplification, Equation 13 and 14 are derived assuming that $\bm{\beta}$ is a deterministic value rather than another distribution. Readers can find more information in [14] for cases where $\bm{\beta}$ also follows a distribution. of the general GPR considering the effect brought by the explicit basis function can then be expressed as444 Equation 13 and 14 assume no noise in observations. If the response has noise that is independently identically distributed and Gaussian $\varepsilon$ with variance $\sigma^{2}$, the kernel function should have one extra term to characterize noise. To this end, the noise kernel function can be written as $\mathbf{K}_{noisy}\left(\mathbf{X},\mathbf{X}\right)=\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)+\sigma^{2}\mathbf{I}$  [^5] Naming convention: $k(\mathbf{x},\mathbf{x})$ represents a number; $\mathbf{k}(\mathbf{x},\mathbf{X})$ represents a $1\times N$ vector and vice verse; $\mathbf{K}(\mathbf{X},\mathbf{X})$ is a $N\times N$ matrix. Moreover, the default vector is a column vector.:.
$\hat{f}(\mathbf{x}^{\prime})=\mathbf{h}(\mathbf{x^{\prime}})^{T}\mathbf{\hat{\boldsymbol{\beta}}}+\mathbf{k}(\mathbf{x}^{\prime},\mathbf{X})\mathbf{K}(\mathbf{X},\mathbf{X})^{-1}(\mathbf{y}-\mathbf{h}(\mathbf{X})^{T}\mathbf{\hat{\boldsymbol{\beta}}})$ (13)
$\hat{\sigma}^{2}(\mathbf{x}^{\prime})=k(\mathbf{x}^{\prime},\mathbf{x}^{\prime})-\mathbf{k}\left(\mathbf{x}^{\prime},\mathbf{X}\right)\mathbf{K}\left(\mathbf{X},\mathbf{X}\right)^{-1}\mathbf{k}\left(\mathbf{X},\mathbf{x}^{\prime}\right)+\mathbf{r}^{T}(\mathbf{h}(\mathbf{X})\mathbf{K}(\mathbf{X},\mathbf{X})^{-1}\mathbf{h}(\mathbf{X})^{T})^{-1}\mathbf{r}$ (14)
where $\mathbf{r}=\mathbf{h}(\mathbf{x^{\prime}})-\mathbf{h}(\mathbf{X})\mathbf{K}(\mathbf{X},\mathbf{X})^{-1}\mathbf{k}(\mathbf{X},\mathbf{x^{\prime}})$. The first two terms are predicted variance deduced by the zero mean GPR and the third term is the predicted variance caused by the explicit basis function.
Incorporating explicit basis functions helps the GPR to generalize better for more complicated problems, while one has to select the basis function $\mathbf{h(x)}$ beforehand or use symbolic regression to identify the terms [73].
@@ A.2.2 Parameters estimation
By taking derivatives of Equation 16 and setting it to zero, we obtain maximum likelihood estimation for $\boldsymbol{\beta}$.
@@
Now the maximum likelihood estimations $\hat{\boldsymbol{\beta}}$ can be substituted back into Equation 16 and constant terms removed to give the concentrated ln-likelihood function:
@@
The value of Equation 18 is determined by parameter $\bm{\theta}$, the maximum value can not be found analytically. Therefore, heuristic algorithms can be utilized to find the optimum value of $\bm{\theta}$.
@@ B.1 LF-KRR parameter estimation
KRR based on the LF dataset can be defined by
@@ B.2 Transfer-learning model and residual GPR parameter estimation
By taking the derivative of Equation 21 and setting it to zero, we can get the estimation of $\hat{\bm{\rho}}$ :
@@ B.3 Predictions
We can obtain the predictions of KRR-LR-GPR by modifying Equation 12 and (14)into:
$\hat{f}^{h}(\mathbf{x}^{\prime})=\mathbf{m}(\mathbf{x^{\prime}})^{T}\mathbf{\hat{\bm{\rho}}}+\mathbf{k}(\mathbf{x}^{\prime},\mathbf{X}^{h})\mathbf{K}(\mathbf{X}^{h},\mathbf{X}^{h})^{-1}(\mathbf{y}^{h}-\mathbf{m}(\mathbf{X}^{h})^{T}\mathbf{\hat{\bm{\rho}}})$ (24)
$\hat{\sigma}_{h}^{2}(\mathbf{x}^{\prime})=k(\mathbf{x}^{\prime},\mathbf{x}^{\prime})-\mathbf{k}\left(\mathbf{x}^{\prime},\mathbf{X}^{h}\right)\mathbf{K}\left(\mathbf{X}^{h},\mathbf{X}^{h}\right)^{-1}\mathbf{k}\left(\mathbf{X}^{h},\mathbf{x}^{\prime}\right)+\mathbf{r}^{T}(\mathbf{m}(\mathbf{X}^{h})\mathbf{K}(\mathbf{X}^{h},\mathbf{X}^{h})^{-1}\mathbf{m}(\mathbf{X}^{h})^{T})^{-1}\mathbf{r}$ (25)
@@ C.1 Numerical functions for testing KRR-LR-GPR
This section includes details on the low-dimensional numerical functions considered in this paper. Originally, those functions were single-fidelity functions used to verify optimization algorithms [74]. They were extended by Jiang et al [75], Sander et al. [76] and Mainini et al. [77] for verifying the performance of MF optimization algorithms.
@@
    where $x\in[0,1]$. By selecting different values of $A$, $B$, and $C$, LF functions with different correlations to the HF function.